# Dyna-Think: Synergizing Reasoning, Acting, and World Model Simulation in AI Agents

### v1

@@ Appendix A Limitations
Benchmark Limitations
In this work, we evaluated on OSWorld (Xie et al., 2024), which to our knowledge is the most diverse computer-use benchmark available. Similar to research in model self-improvement, our approach aims to improve AI agent’s performance mainly using data created by itself (i.e., from online rollouts). In our main experiments, we thus compared different models and methods across 5 domains (OS, VSCode, Chrome, GIMP, and Thunderbird), and excluded domains such as VLC and LibreOffice Impress that frequently encounter representation issues (e.g., VLC interacts with a video file, and LibreOffice Impress requires visual changes to a powerpoint slide). In Section E.1, we find these domains results in poor performance even after distillation from R1. We believe methods that bootstrap model performance from zero would be highly valuable in these domains. We leave this for future work.
Appendix B Ethics Statement
@@
Appendix C Dyna-Think Imitation Learning Details
Figure A1: Example reconstruction in DIT. World model simulation is in green.
@@ Appendix C Dyna-Think Imitation Learning Details
We use GPT-4o to reconstruct R1-style reasoning in Section 3.2. For each turn in an R1-generated trajectory, we prompt GPT-4o to reconstruct $a_{i}$ given its immediate previous observation $o_{i-1}$ and relevant system instructions. We present our prompt in Table A3. A reconstructed output is shown in Figure A1. For R1 responses within 550 tokens, we do not perform abbreviation as these responses already mostly only constitutes verification, world model simulation, and action. To help ensure GPT-4o removes as much token as possible, for each prompt we sample 5 outputs, and use the shortest output in our DIT training corpus.
Appendix D Dyna-Think Dyna Training Details
D.1 World Model Data Prompts
@@
To construct world model data, we experimented with three methods in Section 3.3. Next-state prediction (DDT($\hat{\mathcal{T}}$)) which trains the model to directly predict the next state; State-difference prediction (DDT($\hat{\mathcal{T}}_{\Delta}$)) which trains the model to predict the difference $\Delta(o_{i},o_{i+1})$; and Simulation-critique generation (DDT($\hat{\mathcal{T}}_{\textrm{critic}}$)), which trains the model to generate a critique for simulation in $a_{i}$. To obtain data for DDT($\hat{\mathcal{T}}_{\Delta}$) we prompt GPT-4o using prompts in Table A4. To obtain data for DDT($\hat{\mathcal{T}}_{\textrm{critic}}$), we first prompt GPT-4o to extract the world model simulation in $a_{i}$ corresponding to its final action, and the prompt it to generate a critique using prompts in Table A5. Example next-state, state-difference, and simulation-critique data is shown in Figure A3.
D.2 Training Details
We present an overview of Dyna-Think Dyna Training in Figure A4. Given a set of rollout trajectories, we first perform world model training and then perform policy training. During policy training, the model $\pi_{\mathcal{W}}(\theta)$ is trained to predict the next action $a_{i}$ given its previous context using correct trajectories. During world model training, we experiment with three data formats: DDT($\hat{\mathcal{T}}$) directly trains the model to predict the next state $o_{i+1}$; DDT($\hat{\mathcal{T}}_{\Delta}$) trains the model to predict a natural language description of the state difference $\Delta({o_{i},o_{i+1}})$ generated by prompting GPT-4o (see Table A4; and DDT($\hat{\mathcal{T}}_{\textrm{critic}}$) trains the model to generate a critique for the world model simulation in $a_{i}$ that caused $o_{i+1}$ (see Table A5).
@@ D.2 Training Details
Figure A2: Example simulation critique data. We first prompt GPT-4o to extract the world model simulation (shown in green) corresponding to the final action; and then prompt GPT-4o to generate an critique (shown in red) based on the extracted simulation and the next state (see Table A5). During training, all tokens except for text in red is masked.
@@
To better localize the critique to its corresponding world model simulation, we additionally 1) prompt GPT-4o to inject the critique back to the original response; and 2) only train on the critique by masking out all other tokens (see Figure A2). To inject a critique back into $a_{i}$, we 1) append an id “[[id=x]]” at the end of every sentence in $a_{i}$; and 2) prompt GPT-4o to output an injection location; and 3) inject the critique back into the response based on the output id.
D.3 Synthetic Task Generation
@@
Since world model training only requires rollout trajectories (regardless of their correctness), we experiment with scaling up world model training by using GPT-4o to generate synthetic tasks for rollouts. Specifically, we prompted GPT-4o using the template in Table A6, and generated 200 tasks for each domain. To ensure diversity between tasks, we prompt GPT-4o to directly generate $N=10$ distinct tasks in a single prompt, and repeat this process 20 times for each domain. After this generation process, we manually inspected multiple task and did not find obvious duplicates or unreasonable tasks.
@@
In practice, we notice that many tasks require additional configuration (e.g., setting up email profile for the Thunderbird domain, or downloading an image to edit in GIMP). To accommodate this, we consider a few domain-specific changes. For all synthetic tasks in the thunderbird domain, we use the same email account configurations as used in the original OSWorld dataset. For the GIMP domain, we augmented our prompt template in Table A6 to additionally include an example GIMP task configuration from the OSWorld dataset, and instructed the model to generate new tasks based on the given configuration (see Table A7). We then use the same task configuration used in the prompt, but replace the instruction with our generated task. This process ensures that the agent has the necessary initialization for the task (e.g., at least an image to edit). For all other domains (OS, Chrome, VSCode), we use an empty configuration as they do not require additional resources.
@@
Given these synthetic tasks, we directly perform rollouts using $\pi_{\mathcal{W}}(\theta)$. We then remove trajectories that did not terminate within a maximum of 30 steps, and use all the remaining rollouts to construct world model training data used in Section 5.1. This resulted in a total of 6467 turns available for training, 8x more than the world model training data used in Section 4.3.
Appendix E OSWorld Details
E.1 Additional Domains
@@
In Table A1, we present our model performance on domains not included in our main except (VLC, LibreOffice Impress, LibreOffice Writer, LibreOffice Calc, and Workflow). While we did observe improvement on average after training, we find these are limited (e.g., 6% improvement in VLC only corresponds to completing 1 additional task). Overall, we find solving tasks in these domains often encounters representation issue, such as the need to interact with a video (in VLC) or making visual changes to a powerpoint slide or document (in LibreOffice tasks). These visual information were often missing or mis-represented in the text-modality (accessibility tree) provided by the OSWorld benchmark. Thus, we mainly compare methods on domains such as OS, VSCode, Chrome, GIMP, and Thunderbird, where most tasks are solvable under the text-modality.
E.2 Task Augmentations
@@
As of the date of this work, OSWorld (Xie et al., 2024) is the largest-scale computer-use benchmark available with 369 tasks covering 10 domains. However, as creating evaluators for computer-use tasks is non-trivial, this benchmark mainly serves as testing performances due to its limited size for training. To this end, we follow recent work (Su et al., 2025) and increase the number of computer-use tasks. Specifically, we manually create new tasks by modifying the original instruction; and/or the task initialization configuration; and the corresponding evaluator configuration. We ensure the resulting task is 1) distinct from the original task, such that action sequences that can correctly complete the original task does not complete the augmented tasks; and 2) can be correctly evaluated by modifying the existing evaluator configurations. We present some example augmented tasks in Table A2.
Appendix F Analysis Details
F.1 World Model Accuracy Test
Given a trajectory from $\pi_{\mathcal{W}}(\theta)$ that performs world model simulation during its thinking process, we 1) first prompt GPT-4o to extract the world model simulation corresponding to the final action $a_{i}$; and then 2) prompt GPT-4o to judge the correctness of this simulation given the next state. We repeat this for every $a_{i}$ in a all trajectories that terminated within a maximum budget of 30 steps. We present the prompts used to extract world model simulation in Table A8, and prompts used to measure simulation accuracy in Table A9.
@@ F.1 World Model Accuracy Test
Table A1: Model performance on additional domains (VLC, LibreOffice Impress, LibreOffice Writer, LibreOffice Calc, and Workflow). All model training are based on Qwen-2.5-32B-Instruct. We excluded these domains from Section 4 because they often require visual interactions (e.g., VLC involves video controls and LibreOffice Impress often needs slide edits) that the accessibility tree often cannot represent.

### v3

@@ Appendix B Limitations
Appendix C Ethics Statement
@@
Appendix D Dyna-Think Imitation Learning Details
Figure A1: Example reconstruction in DIT. World simulation is in green.
@@ Appendix D Dyna-Think Imitation Learning Details
We use GPT-4o to reconstruct R1-style reasoning in Section˜3.2. For each turn in an R1-generated trajectory, we prompt GPT-4o to reconstruct $a_{i}$ given its immediate previous observation $o_{i-1}$ and relevant system instructions. We present our prompt in Table˜A4. A reconstructed output is shown in Figure˜A1. For R1 responses within 550 tokens, we do not perform abbreviation as these responses already mostly only constitutes verification, world model simulation, and action. To help ensure GPT-4o removes as much token as possible, for each prompt we sample 5 outputs, and use the shortest output in our DIT training corpus.
Appendix E Dyna-Think Dyna Training Details
E.1 World Model Data Prompts
@@
To construct world model data, we experimented with three methods in Section˜3.3. Next-state prediction (DDT($\hat{\mathcal{T}}$)) which trains the model to directly predict the next state; State-difference prediction (DDT($\hat{\mathcal{T}}_{\Delta}$)) which trains the model to predict the difference $\Delta(o_{i},o_{i+1})$; and Simulation-critique generation (DDT($\hat{\mathcal{T}}_{\textrm{critic}}$)), which trains the model to generate a critique for simulation in $a_{i}$. To obtain data for DDT($\hat{\mathcal{T}}_{\Delta}$) we prompt GPT-4o using prompts in Table˜A5. To obtain data for DDT($\hat{\mathcal{T}}_{\textrm{critic}}$), we first prompt GPT-4o to extract the world model simulation in $a_{i}$ corresponding to its final action, and the prompt it to generate a critique using prompts in Table˜A6. Example next-state, state-difference, and simulation-critique data is shown in Figure˜A3.
E.2 Training Details
@@
Figure A2: Example simulation critique data. We first prompt GPT-4o to extract the world model simulation (shown in green) corresponding to the final action; and then prompt GPT-4o to generate an critique (shown in red) based on the extracted simulation and the next state (see Table˜A6). During training, all tokens except for text in red is masked.
We present an overview of Dyna-Think Dyna Training in Figure˜A4. Given a set of rollout trajectories, we first perform world model training and then perform policy training. During policy training, the model $\pi_{\mathcal{W}}(\theta)$ is trained to predict the next action $a_{i}$ given its previous context using correct trajectories. During world model training, we experiment with three data formats: DDT($\hat{\mathcal{T}}$) directly trains the model to predict the next state $o_{i+1}$; DDT($\hat{\mathcal{T}}_{\Delta}$) trains the model to predict a natural language description of the state difference $\Delta({o_{i},o_{i+1}})$ generated by prompting GPT-4o (see Table˜A5; and DDT($\hat{\mathcal{T}}_{\textrm{critic}}$) trains the model to generate a critique for the world model simulation in $a_{i}$ that caused $o_{i+1}$ (see Table˜A6).
@@
To better localize the critique to its corresponding world model simulation, we additionally 1) prompt GPT-4o to inject the critique back to the original response; and 2) only train on the critique by masking out all other tokens (see Figure˜A2). To inject a critique back into $a_{i}$, we 1) append an id “[[id=x]]” at the end of every sentence in $a_{i}$; and 2) prompt GPT-4o to output an injection location; and 3) inject the critique back into the response based on the output id.
In Algorithm˜1 we provide a high-level pseudocode for DDT. For world model training, we adopt Language Modeling(LM) Loss:
$\mathcal{L}_{\text{LM}}(\theta)=-\sum_{t=1}^{T}\log\pi(x_{t}\mid x_{<t}),$ (1)
During policy training, we use policy gradient to optimize the reward:
$\nabla_{\theta}\mathcal{L}_{\text{policy}}(\theta)=\mathbb{E}_{\tau\sim\pi_{\mathcal{W}}(\theta)}\Big[\sum_{t=0}^{T}\nabla_{\theta}\log\pi(a_{t}\mid o_{t})R(\tau)\Big],$ (2)
where we use $R(\tau)=1$ for successful tasks, otherwise 0. In our main experiment (Table˜1), we perform $N=1$ iteration enable to have a fair comparison between the data used for policy and world model training across different algorithms (RFT and vanilla Dyna). In Section˜5.2 we extend our method to multiple iterations.
Algorithm 1 DDT Training (Policy + World Model with Critique Injection)
1:Policy/world model $\pi_{W}(\theta)$, environment $\mathcal{T}$, rollout trajectory $\tau$, reward function $R$
2:repeat NN times $\triangleright$ Data Collection:
3:  Roll out $\pi_{\mathcal{W}}(\theta)$ in $\mathcal{T}$ for trajectory $\tau=(o_{0},a_{1},o_{1},a_{2},\ldots,a_{T})$
4:  Construct policy dataset $\mathcal{D}_{\pi}\leftarrow\{(\text{context}(o_{0},a_{<i},o_{<i}),a_{i},R)\}$
5:  Construct world model dataset $\mathcal{D}_{W}\leftarrow\{(o_{i},a_{i},o_{i+1})\}$ $\triangleright$ World Model Training:
6:  for all $(o_{i},a_{i},o_{i+1})\in\mathcal{D}_{W}$ do
7:   Option 1: Next-state prediction (DDT($\hat{\mathcal{T}}$))
8:     Input $(o_{i},a_{i})$, target $o_{i+1}$
9:     $\mathcal{L}_{\text{wm}}\leftarrow\mathcal{L}_{\text{LM}}(\pi_{W}(\theta),(o_{i},a_{i}),o_{i+1})$
10:   Option 2: State-change prediction (DDT($\hat{\mathcal{T}}_{\Delta}$))
11:     Input $(o_{i},a_{i})$, target $\Delta(o_{i},o_{i+1})$ (NL description from GPT-4o)
12:     $\mathcal{L}_{\text{wm}}\leftarrow\mathcal{L}_{\text{LM}}(\pi_{W}(\theta),(o_{i},a_{i}),\Delta(o_{i},o_{i+1}))$
13:   Option 3: Critique prediction (DDT($\hat{\mathcal{T}}_{\text{critic}}$))
14:     (1) Append IDs [[id=x]] to each sentence in $a_{i}$
15:     (2) Use GPT-4o to output injection location
16:     (3) Inject critique $\to a^{\prime}_{i}$
17:     (4) Mask non-critique tokens in $a^{\prime}_{i}$
18:     $\mathcal{L}_{\text{wm}}\leftarrow\mathcal{L}_{\text{LM}}(\pi_{W}(\theta),(o_{i},\ldots),\text{masked critique tokens})$
19:  end for
20:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}L_{\text{wm}}$ $\triangleright$ Policy Training:
21:  for all $(\text{context},a_{i},R)\in\mathcal{D}_{\pi}$ do
22:   $\mathcal{L}_{\pi}\leftarrow\mathcal{L}_{\text{policy}}(\pi_{W}(\theta),\text{context},a_{i};R)$
23:  end for
24:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\pi}$
25:end
26:return $\pi_{\mathcal{W}}(\theta)$
E.3 Synthetic Task Generation
@@
Since world model training only requires rollout trajectories (regardless of their correctness), we experiment with scaling up world model training by using GPT-4o to generate synthetic tasks for rollouts. Specifically, we prompted GPT-4o using the template in Table˜A7, and generated 200 tasks for each domain. To ensure diversity between tasks, we prompt GPT-4o to directly generate $N=10$ distinct tasks in a single prompt, and repeat this process 20 times for each domain. After this generation process, we manually inspected multiple task and did not find obvious duplicates or unreasonable tasks.
@@
In practice, we notice that many tasks require additional configuration (e.g., setting up email profile for the Thunderbird domain, or downloading an image to edit in GIMP). To accommodate this, we consider a few domain-specific changes. For all synthetic tasks in the thunderbird domain, we use the same email account configurations as used in the original OSWorld dataset. For the GIMP domain, we augmented our prompt template in Table˜A7 to additionally include an example GIMP task configuration from the OSWorld dataset, and instructed the model to generate new tasks based on the given configuration (see Table˜A8). We then use the same task configuration used in the prompt, but replace the instruction with our generated task. This process ensures that the agent has the necessary initialization for the task (e.g., at least an image to edit). For all other domains (OS, Chrome, VSCode), we use an empty configuration as they do not require additional resources.
@@
Given these synthetic tasks, we directly perform rollouts using $\pi_{\mathcal{W}}(\theta)$. We then remove trajectories that did not terminate within a maximum of 30 steps, and use all the remaining rollouts to construct world model training data used in Section˜5.1. This resulted in a total of 6467 turns available for training, 8x more than the world model training data used in Section˜4.3.
Appendix F OSWorld Details
F.1 Additional Domains
@@
In Table˜A2, we present our model performance on domains not included in our main except (VLC, LibreOffice Impress, LibreOffice Writer, LibreOffice Calc, and Workflow). While we did observe improvement on average after training, we find these are limited (e.g., 6% improvement in VLC only corresponds to completing 1 additional task). Overall, we find solving tasks in these domains often encounters representation issue, such as the need to interact with a video (in VLC) or making visual changes to a powerpoint slide or document (in LibreOffice tasks). These visual information were often missing or mis-represented in the text-modality (accessibility tree) provided by the OSWorld benchmark. Thus, we mainly compare methods on domains such as OS, VSCode, Chrome, GIMP, and Thunderbird, where most tasks are solvable under the text-modality.
F.2 Task Augmentations
@@
Table A1: Number of tasks per domain in the training and test set.
Domain Train Test
OS 57 31
VSCode 46 38
Chrome 88 54
GIMP - 32
TBird - 19
As of the date of this work, OSWorld (Xie et al., 2024) is the largest-scale computer-use benchmark available with 369 tasks covering 10 domains. However, as creating evaluators for computer-use tasks is non-trivial, this benchmark mainly serves as testing performances due to its limited size for training. To this end, we follow recent work (Su et al., 2025) and increase the number of computer-use tasks. Specifically, we manually create new tasks by modifying the original instruction; and/or the task initialization configuration; and the corresponding evaluator configuration. We ensure the resulting task is 1) distinct from the original task, such that action sequences that can correctly complete the original task does not complete the augmented tasks; and 2) can be correctly evaluated by modifying the existing evaluator configurations. We present some example augmented tasks in Table˜A3.
Appendix G Analysis Details
G.1 World Model Accuracy Test
Given a trajectory from $\pi_{\mathcal{W}}(\theta)$ that performs world model simulation during its thinking process, we 1) first prompt GPT-4o to extract the world model simulation corresponding to the final action $a_{i}$; and then 2) prompt GPT-4o to judge the correctness of this simulation given the next state. We repeat this for every $a_{i}$ in a all trajectories that terminated within a maximum budget of 30 steps. We present the prompts used to extract world model simulation in Table˜A9, and prompts used to measure simulation accuracy in Table˜A10.
@@ G.1 World Model Accuracy Test
Table A2: Model performance on additional domains (VLC, LibreOffice Impress, LibreOffice Writer, LibreOffice Calc, and Workflow). All model training are based on Qwen-2.5-32B-Instruct. We excluded these domains from Section˜4 because they often require visual interactions (e.g., VLC involves video controls and LibreOffice Impress often needs slide edits) that the accessibility tree often cannot represent.